3. Suppose x,y,z satisfy the competing species equations <(6 - 2x – 3y - 2) y(7 - 2x - 3y - 22) z(5 - 2x - y -22) (a) (6 poin(c) (8 points) Determine if the critical point (2,0,0) is stable, where I > 0.
Answer #1

y=0, 3y + 2z=7 and The y nullelines aree z-null elines aree, z=0, Y + 2z=5. Freom the above figure, that Z-ncellclines and ynfigure of nullclines in yz-plane, is drawn below

2.U 39,732 = 0 9:10+ = -0. 9. 0 8. E O 7. lo 5- 4. 3. 2+. 1.0 1. 0.0 0.0 0.0 1 2 2.53 3.5 5 6 X -1 -2-Y Now put y=0, Z=0 in system 11, then CD changes to dx x 16-2x). Now, dx de =0 > x=0, x=3 The Critical point is 13,0,0) wherc